# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v2

@@ 3.4 Consensus & Optimistic Approaches
Verde [41] employs optimistic dispute resolution, recomputing only divergent operators (e.g., single attention heads) to minimize overhead. Mira [42] leverages multi-model consensus against hallucinations using domain-specific thresholds. Ambient [43] introduces Proof-of-Logits (PoL), hashing intermediate logits for spot-check verification at $0.1\%$ overhead. However, they introduce delayed finality during dispute windows and implicitly assume honest majority participation.
@@ 4 Background & Problem Statements
Full-Sequence Prefill for Parallel Verification. Autoregressive large language models (LLMs) rely on causal (left-to-right) self-attention, in which each hidden state $h_{t}$ depends on all preceding tokens $\{x_{1},\dots,x_{t-1}\}$. This dependency constrains the Decoder to operate sequentially, forming the dominant computational bottleneck in inference. Practical serving systems mitigate this through a Prefill phase that computes the contextual representations for all prompt tokens in parallel and materializes the key–value (KV) cache, allowing subsequent token-by-token decoding to reuse these cached features [48, 68, 51]. At verification time, this architectural separation becomes even more advantageous: since the verifier observes the full input–output sequence $(x_{1},\dots,x_{T})$, it can execute a single forward pass under the causal mask to recover all hidden states simultaneously, without performing the autoregressive loop.
@@
Lazy Verification. A naive verification baseline re-executes the entire model forward pass under identical hyperparameters and compares the resulting hidden states with those reported by the Inferencer. However, because honest inference is the common case, verifiers have an incentive to skip computation and submit a trivial True verdict—an instance of the lazy-verification problem. Such behavior gradually undermines system integrity and leads to unverifiable results.
To address this, VeriLLM employs a lightweight commit–sample–check protocol. Each verifier $v_{i}$ commits to its hidden states by posting a Merkle root $r_{i}=\mathsf{Merkle}(H_{i})$ [44]. After commitments are finalized, a smart contract derives random challenge indices $\mathcal{S}=\{s_{1},\dots,s_{k}\}$ from a verifiable random function (VRF) [45], ensuring unpredictability and fairness. Each verifier must open the corresponding leaves $\{H_{i}[s_{j}]\}_{j=1}^{k}$ and submit inclusion proofs $\pi_{i}[s_{j}]$ to the contract. Let $\delta(H_{i}[s_{j}],H^{*}[s_{j}])$ denote the numerical deviation from the Inferencer’s committed states. The contract enforces:
@@
where $\epsilon$ is a tolerance threshold capturing floating-point nondeterminism [69, 70]. Verifiers whose openings align with the Inferencer’s states within $\epsilon$ receive rewards, while inconsistent openings trigger slashing. To ensure continuous participation, a peer-consistency mechanism statistically compares verifier outputs and rewards correlation with the empirical distribution expected from honest computation, thus discouraging collusion and inactive nodes.
Security under Sampling Majority. Conventional consensus protocols [25, 26] rely on an honest-majority assumption to guarantee safety and liveness. In contrast, VeriLLM achieves correctness under the weaker one-honest-verifier (OHV) assumption. Let $n$ denote the total number of verifiers, $f$ the number of malicious ones, and $k$ the committee size for sampling-based verification. The probability that all $k$ selected verifiers are malicious is:
@@ 5.1 System Overview
  • Scheduler. The scheduler orchestrates task placement, resource allocation, and verification planning. For each inference job, it selects a candidate set of node groups capable of jointly executing the model’s layers or tensor partitions. Within each selected group, the scheduler designates one Inferencer and several Verifiers to handle computation and validation, respectively. Although logically centralized, the scheduler’s randomness and decisions are publicly verifiable: both role assignment and checkpoint sampling derive from a verifiable random function (VRF) [45], producing unpredictable yet reproducible outcomes. This prevents adversaries from precomputing which nodes will be responsible for verification or from biasing verifier selection. The scheduler also records the mapping of job identifiers, selected groups, and Merkle root commitments of intermediate states onto the blockchain for immutable traceability.
@@
  • Node Groups. VeriLLM organizes computation across groups of homogeneous GPU workers to simplify performance balancing and state consistency. Each node group corresponds to one or more identical Transformer layers, including all weights, normalization statistics, and kernel parameters. This design aligns with modern distributed inference frameworks such as tensor- and pipeline-parallel architectures [51, 52]. Within a group, a node can serve either as an Inferencer or a Verifier depending on the scheduler’s role assignment for the current job. To mitigate role-conditioned adversarial strategies, both roles share a unified execution interface: from the node’s perspective, the input tensor, KV-cache access pattern, and runtime sequence length appear identical for inference and verification. Consequently, at the start of the Prefill phase, a worker cannot distinguish whether it is performing live inference or verifying another node’s computation. This task-type indistinguishability provides a built-in security property that limits strategic deviation, since any deviation risks immediate detection if the task later turns out to be a verification probe.
@@
  • Verification Contracts. The verification layer is implemented through on-chain smart contracts [71, 72]. These contracts collect verifier submissions, validate sampling correctness, and finalize verification outcomes through contract-level consensus. Each participating node is required to deposit a stake when registering for inference or verification. The contracts automatically enforce rewards and penalties based on the outcome of verification, slashing misbehaving nodes when inconsistencies or invalid openings are detected. To minimize gas costs, VeriLLM adopts a two-tier protocol: (i) the heavy-lift computation and Merkle proof generation occur off-chain; (ii) only small digests, VRF proofs, and sampled Merkle openings are submitted on-chain. This hybrid design maintains full public verifiability while keeping on-chain cost sublinear in model size. Moreover, the contract interface exposes an event log of all verification transactions, allowing any third-party observer to independently re-verify computations, thereby realizing trustless auditability.
@@ 5.2 Registration
    the scheduler’s VRF verification public key, enabling third parties to validate the pseudorandom outputs used in node assignment and hidden-state sampling [45];
@@
The VRF key ensures that all sampling and scheduling randomness remains publicly auditable and resistant to precomputation or selective bias [71, 72].
@@
This mechanism parallels the delayed-withdrawal model of accountable distributed systems [25, 73], guaranteeing that both schedulers and workers remain economically bound to their past behavior.
@@ 5.3 Scheduling Scheme
Verifiable randomness and auditability. Each scheduling decision is deterministically derived from a publicly verifiable random beacon instantiated via a verifiable random function (VRF) [45]. The Scheduler’s VRF verification key is registered on-chain, and every scheduling round includes the tuple:
@@ 5.4 Task Scheduling and Hidden-State Management
ensuring accountability for both the producer and the forwarder. This dual-signature commitment model provides end-to-end traceability: any discrepancy between a segment’s output and the forwarded record can be attributed with cryptographic certainty. Because verifiers only require Merkle openings for sampled positions, the bandwidth and on-chain verification costs scale sublinearly with tensor size. The integration of deterministic serialization, Merkle-based commitment [44], and chained signatures provides tamper-evident provenance for all hidden states without introducing measurable latency into the inference path.
Overall, this architecture ensures that every stage of distributed inference remains cryptographically auditable while preserving throughput and latency comparable to optimized large-model serving systems [51, 52].
@@ 5.5 Inference Scheme
The inference process executes online, autoregressive generation over a decentralized, pipelined fabric of homogeneous node groups. Each group hosts an identical copy of a contiguous Transformer block, including weights, normalization statistics, and kernel configurations [52, 51]. For each incoming task, the Scheduler selects eligible groups and, through a verifiable random function (VRF) [45], assigns one Inferencer per group to execute the live forward pass. Verifier roles are simultaneously drawn from the same VRF distribution but remain inactive during this phase. Throughout generation, the Scheduler relays and logs hidden states across segment boundaries, appending cryptographic commitments and signatures to construct an auditable trace that serves as the foundation for subsequent verification [27, 38].
@@
where $S^{(t)}_{i}$ is the hidden state entering segment $i$, $F_{i}$ denotes the segment’s forward operator, $W_{i}$ its parameters, and Decode the final output head combined with the sampling rule. The resulting pipeline achieves token-serial yet segment-parallel execution, synchronized by the Scheduler’s relay of hidden states [51].
Scheduler-Driven Role Assignment. For each inference task, the Scheduler enumerates all node groups capable of collectively serving the model and assigns roles in a publicly verifiable manner. Using a VRF bound to the task identifier [45], it deterministically selects one Inferencer per group, while keeping the randomness unpredictable before the VRF proof is revealed. This mechanism ensures that role assignment cannot be biased or precomputed and that every selection remains verifiable through the VRF proof posted on-chain [50, 40].
@@
    computes a cryptographic commitment to $S^{(t)}_{i+1}$ (e.g., a Merkle root over a deterministic serialization [44]) and signs the root with its node key; and
@@
The Scheduler verifies each signature against the on-chain registry (optionally through a zero-knowledge membership proof [38, 39] to preserve role anonymity), appends its own relay signature, persists the state to stable storage, and forwards the package to the next segment’s Inferencer. To minimize latency, each receiving node can begin its computation immediately while performing signature verification asynchronously; if any proof later fails, the Scheduler aborts the affected pipeline and invalidates dependent results.
@@
All node public keys are registered under an on-chain Merkle accumulator, enabling compact membership proofs [45, 50]. Instead of forwarding raw signatures, the Scheduler may alternatively issue a zero-knowledge proof attesting that it holds a valid signature from some registered key on the advertised root—thereby maintaining verifiability without revealing node identity [39, 40]. Each forwarded artifact is re-signed by the Scheduler to guarantee origin authenticity and replay protection. This layered commitment and signature discipline ensures that every hop in the relay chain is both tamper-evident and cryptographically attributable.
Autoregressive Token Emission and Feedback. The terminal segment ($i=L$) outputs logits or the final hidden state for token $t$, from which the next token $y_{t}$ is sampled according to the decoding policy [2]. The Scheduler collects $y_{t}$ and feeds it back into the pipeline to form $x_{1:t+1}$, advancing the process to step t+1t{+}1. The loop continues until the model emits an end-of-sequence token $\langle\texttt{EOS}\rangle$ or another termination condition is met.
Task-Type Indistinguishability and Robustness. The communication protocol, message format, and API surface used by Inferencers during inference are identical to those used by Verifiers in the subsequent verification phase [11]. Consequently, no node can distinguish in real time whether it is performing live inference or participating in a verification probe. This indistinguishability eliminates the opportunity for strategic deviation and simplifies implementation symmetry across both phases. Combined with VRF-driven role assignment, signed commitments, and auditable relays, the inference scheme produces a verifiable, tamper-evident execution trace with negligible online overhead.
@@ 5.6 Verification Scheme
During autoregressive inference, an LLM executes two phases: a prefill pass—where the prompt tokens are processed in parallel to populate the key–value (KV) cache—and a decode phase—where tokens are generated sequentially with causal dependencies on all previous outputs [2, 52]. The decoding phase dominates latency and is inherently serial, leading to limited GPU utilization and memory-bound performance [51]. However, verification differs fundamentally: once the full output sequence $y_{1:T}$ is known, the verifier is no longer constrained by sequential dependencies. Instead, it can recompute all token representations simultaneously under the same causal mask, yielding a single parallel pass equivalent to prefill over the concatenated prompt–output sequence [39, 27].
@@
and let $\mathcal{S}=[S^{\text{prompt}}\|S^{1}\|\dots\|S^{T}]$ denote the concatenated hidden states across all positions. Given the causal structure of the Transformer, the verifier can execute a single forward pass to reconstruct all hidden states $\{S^{(t)}_{i}\}_{t=1}^{T}$ for every segment $i\in[1,L]$ without performing incremental token decoding [74]. This transformation replaces $T$ serial decode iterations with a single batched forward computation, eliminating redundant KV-cache updates and reducing cross-node communication. In practice, this yields significantly improved GPU utilization, especially on architectures with high Tensor Core throughput [51].
@@
yielding near-linear scaling in the number of participating verifier groups [51, 27].
@@
Prefill Computation. After the inference phase concludes, the Scheduler initiates segment-local verification by distributing the required input states to each verifier [40]. For segment $i>1$, the Scheduler sends the sequence $\{S_{i}^{(t)}\}_{t=1}^{T}$—the hidden states produced by the upstream inferencer for all tokens. For the first segment ($i=1$), the Scheduler instead provides the full tokenized input $\texttt{prompt}\|y_{1:T}]$, corresponding to the raw embedding input of the model. Upon receipt, each verifier executes a full-sequence prefill limited to its hosted layers:
@@
Importantly, the communication interface, data serialization format, and RPC protocol used during verification are identical to those used in live inference [11]. This design choice enforces task-type indistinguishability: from the perspective of any node, it is impossible to discern whether the current request corresponds to live inference or a verification probe [50]. Such indistinguishability eliminates mode-specific behavior and prevents nodes from selectively deviating or shortcutting computation when acting as verifiers. Upon completion, each verifier must return its computed outputs $\{S^{(t)}_{i+1}\}_{t=1}^{T}$ through the same interface as during inference, producing an auditable trace indistinguishable in structure and timing.
Commitment Submission. Once the local recomputation completes, the Scheduler transitions verifiers into the commit phase. Each verifier constructs a tamper-evident cryptographic commitment to its final outputs—specifically, the hidden state corresponding to the last token ($t=T$) at its segment’s output boundary. This commitment is instantiated as a Merkle root [44]:
@@
and publishes the commitment on-chain through the designated verification contract [38, 39].
This on-chain record provides a permanent anchor for later verification steps. In subsequent sampling or dispute phases, verifiers may be required to disclose specific slices or elements of the hidden state $\{S^{(T)}_{i+1}\}$, accompanied by Merkle inclusion proofs that can be efficiently checked against the posted root [50, 45]. This approach ensures both compactness and auditability: the entire verification process involves only lightweight commitments and selective disclosures rather than full tensor uploads, keeping blockchain overhead minimal while preserving cryptographic integrity.
In summary, the verification scheme leverages the known output sequence and the pre-recorded intermediate states to convert sequential autoregressive computation into parallel, segment-local prefill verification. Combined with Merkle commitments and signature anchoring, this design achieves strong verifiability and high throughput, ensuring that even a single honest verifier can detect and prove any inference deviation with overwhelming probability [40, 39].

### v3

@@ 3.4 Consensus & Optimistic Approaches
Verde [arun2025verde] employs optimistic dispute resolution, recomputing only divergent operators (e.g., single attention heads) to minimize overhead. Mira [mira-network] leverages multi-model consensus against hallucinations using domain-specific thresholds. Ambient [ambient-ai] introduces Proof-of-Logits (PoL), hashing intermediate logits for spot-check verification at $0.1\%$ overhead. However, they introduce delayed finality during dispute windows and implicitly assume honest majority participation.
@@ 4 Background & Problem Statements
Full-Sequence Prefill for Parallel Verification. Autoregressive large language models (LLMs) rely on causal (left-to-right) self-attention, in which each hidden state $h_{t}$ depends on all preceding tokens $\{x_{1},\dots,x_{t-1}\}$. This dependency constrains the Decoder to operate sequentially, forming the dominant computational bottleneck in inference. Practical serving systems mitigate this through a Prefill phase that computes the contextual representations for all prompt tokens in parallel and materializes the key–value (KV) cache, allowing subsequent token-by-token decoding to reuse these cached features [agrawal2023, shazeer2019fast, rajbhandari2022deep]. At verification time, this architectural separation becomes even more advantageous: since the verifier observes the full input–output sequence $(x_{1},\dots,x_{T})$, it can execute a single forward pass under the causal mask to recover all hidden states simultaneously, without performing the autoregressive loop.
@@
Lazy Verification. A naive verification baseline re-executes the entire model forward pass under identical hyperparameters and compares the resulting hidden states with those reported by the Inferencer. However, because honest inference is the common case, verifiers have an incentive to skip computation and submit a trivial True verdict—an instance of the lazy-verification problem, aka. the verifier’s dilemma [luu2015demystifying]. Such behavior gradually undermines system integrity and leads to unverifiable results.
To address this, VeriLLM employs a lightweight commit–sample–check protocol. Each verifier $v_{i}$ commits to its hidden states by posting a Merkle root $r_{i}=\mathsf{Merkle}(H_{i})$ [merkle1989certified]. After commitments are finalized, a smart contract derives random challenge indices $\mathcal{S}=\{s_{1},\dots,s_{k}\}$ from a verifiable random function (VRF) [micali1999verifiable], ensuring unpredictability and fairness. Each verifier must open the corresponding leaves $\{H_{i}[s_{j}]\}_{j=1}^{k}$ and submit inclusion proofs $\pi_{i}[s_{j}]$ to the contract. Let $\delta(H_{i}[s_{j}],H^{*}[s_{j}])$ denote the numerical deviation from the Inferencer’s committed states. The contract enforces:
@@
where $\epsilon$ is a tolerance threshold capturing floating-point nondeterminism [dettmers2022llmint8, frantar2023gptq]. Verifiers whose openings align with the Inferencer’s states within $\epsilon$ receive rewards, while inconsistent openings trigger slashing. To ensure continuous participation, a peer-consistency mechanism statistically compares verifier outputs and rewards correlation with the empirical distribution expected from honest computation, thus discouraging collusion and inactive nodes.
Security under Sampling Majority. Conventional consensus protocols [castro1999practical, cachin2011introduction] rely on an honest-majority assumption to guarantee safety and liveness. In contrast, VeriLLM achieves correctness under the weaker one-honest-verifier (OHV) assumption. Let $n$ denote the total number of verifiers, $f$ the number of malicious ones, and $k$ the committee size for sampling-based verification. The probability that all $k$ selected verifiers are malicious is:
@@ 5.1 System Overview
  • Scheduler. The scheduler orchestrates task placement, resource allocation, and verification planning. For each inference job, it selects a candidate set of node groups capable of jointly executing the model’s layers or tensor partitions. Within each selected group, the scheduler designates one Inferencer and several Verifiers to handle computation and validation, respectively. Although logically centralized, the scheduler’s randomness and decisions are publicly verifiable: both role assignment and checkpoint sampling derive from a verifiable random function (VRF) [micali1999verifiable], producing unpredictable yet reproducible outcomes. This prevents adversaries from precomputing which nodes will be responsible for verification or from biasing verifier selection. The scheduler also records the mapping of job identifiers, selected groups, and Merkle root commitments of intermediate states onto the blockchain for immutable traceability.
@@
  • Node Groups. VeriLLM organizes computation across groups of homogeneous GPU workers to simplify performance balancing and state consistency. Each node group corresponds to one or more identical Transformer layers, including all weights, normalization statistics, and kernel parameters. This design aligns with modern distributed inference frameworks such as tensor- and pipeline-parallel architectures [rajbhandari2022deep, shoeybi2019megatron]. Within a group, a node can serve either as an Inferencer or a Verifier depending on the scheduler’s role assignment for the current job. To mitigate role-conditioned adversarial strategies, both roles share a unified execution interface: from the node’s perspective, the input tensor, KV-cache access pattern, and runtime sequence length appear identical for inference and verification. Consequently, at the start of the Prefill phase, a worker cannot distinguish whether it is performing live inference or verifying another node’s computation. This task-type indistinguishability provides a built-in security property that limits strategic deviation, since any deviation risks immediate detection if the task later turns out to be a verification probe.
@@
  • Verification Contracts. The verification layer is implemented through on-chain smart contracts [buterin2014next, wood2014ethereum]. These contracts collect verifier submissions, validate sampling correctness, and finalize verification outcomes through contract-level consensus. Each participating node is required to deposit a stake when registering for inference or verification. The contracts automatically enforce rewards and penalties based on the outcome of verification, slashing misbehaving nodes when inconsistencies or invalid openings are detected. To minimize gas costs, VeriLLM adopts a two-tier protocol: (i) the heavy-lift computation and Merkle proof generation occur off-chain; (ii) only small digests, VRF proofs, and sampled Merkle openings are submitted on-chain. This hybrid design maintains full public verifiability while keeping on-chain cost sublinear in model size. Moreover, the contract interface exposes an event log of all verification transactions, allowing any third-party observer to independently re-verify computations, thereby realizing trustless auditability.
@@ 5.2 Registration
    the scheduler’s VRF verification public key, enabling third parties to validate the pseudorandom outputs used in node assignment and hidden-state sampling [micali1999verifiable];
@@
The VRF key ensures that all sampling and scheduling randomness remains publicly auditable and resistant to precomputation or selective bias [buterin2014next, wood2014ethereum].
@@
This mechanism parallels the delayed-withdrawal model of accountable distributed systems [castro1999practical, luu2016smartpool], guaranteeing that both schedulers and workers remain economically bound to their past behavior.
@@ 5.3 Scheduling Scheme
Verifiable randomness and auditability. Each scheduling decision is deterministically derived from a publicly verifiable random beacon instantiated via a verifiable random function (VRF) [micali1999verifiable]. The Scheduler’s VRF verification key is registered on-chain, and every scheduling round includes the tuple:
@@ 5.4 Task Scheduling and Hidden-State Management
ensuring accountability for both the producer and the forwarder. This dual-signature commitment model provides end-to-end traceability: any discrepancy between a segment’s output and the forwarded record can be attributed with cryptographic certainty. Because verifiers only require Merkle openings for sampled positions, the bandwidth and on-chain verification costs scale sublinearly with tensor size. The integration of deterministic serialization, Merkle-based commitment [merkle1989certified], and chained signatures provides tamper-evident provenance for all hidden states without introducing measurable latency into the inference path.
Overall, this architecture ensures that every stage of distributed inference remains cryptographically auditable while preserving throughput and latency comparable to optimized large-model serving systems [rajbhandari2022deep, shoeybi2019megatron].
@@ 5.5 Inference Scheme
The inference process executes online, autoregressive generation over a decentralized, pipelined fabric of homogeneous node groups. Each group hosts an identical copy of a contiguous Transformer block, including weights, normalization statistics, and kernel configurations [shoeybi2019megatron, rajbhandari2022deep]. For each incoming task, the Scheduler selects eligible groups and, through a verifiable random function (VRF) [micali1999verifiable], assigns one Inferencer per group to execute the live forward pass. Verifier roles are simultaneously drawn from the same VRF distribution but remain inactive during this phase. Throughout generation, the Scheduler relays and logs hidden states across segment boundaries, appending cryptographic commitments and signatures to construct an auditable trace that serves as the foundation for subsequent verification [zhang2020delphi, chen2024zkml].
@@
where $S^{(t)}_{i}$ is the hidden state entering segment $i$, $F_{i}$ denotes the segment’s forward operator, $W_{i}$ its parameters, and Decode the final output head combined with the sampling rule. The resulting pipeline achieves token-serial yet segment-parallel execution, synchronized by the Scheduler’s relay of hidden states [rajbhandari2022deep].
Scheduler-Driven Role Assignment. For each inference task, the Scheduler enumerates all node groups capable of collectively serving the model and assigns roles in a publicly verifiable manner. Using a VRF bound to the task identifier [micali1999verifiable], it deterministically selects one Inferencer per group, while keeping the randomness unpredictable before the VRF proof is revealed. This mechanism ensures that role assignment cannot be biased or precomputed and that every selection remains verifiable through the VRF proof posted on-chain [bonneau2015sok, qu2025zkgpt].
@@
    computes a cryptographic commitment to $S^{(t)}_{i+1}$ (e.g., a Merkle root over a deterministic serialization [merkle1989certified]) and signs the root with its node key; and
@@
The Scheduler verifies each signature against the on-chain registry (optionally through a zero-knowledge membership proof [chen2024zkml, sun2024zkllm] to preserve role anonymity), appends its own relay signature, persists the state to stable storage, and forwards the package to the next segment’s Inferencer. To minimize latency, each receiving node can begin its computation immediately while performing signature verification asynchronously; if any proof later fails, the Scheduler aborts the affected pipeline and invalidates dependent results.
@@
All node public keys are registered under an on-chain Merkle accumulator, enabling compact membership proofs [micali1999verifiable, bonneau2015sok]. Instead of forwarding raw signatures, the Scheduler may alternatively issue a zero-knowledge proof attesting that it holds a valid signature from some registered key on the advertised root—thereby maintaining verifiability without revealing node identity [sun2024zkllm, qu2025zkgpt]. Each forwarded artifact is re-signed by the Scheduler to guarantee origin authenticity and replay protection. This layered commitment and signature discipline ensures that every hop in the relay chain is both tamper-evident and cryptographically attributable.
Autoregressive Token Emission and Feedback. The terminal segment ($i=L$) outputs logits or the final hidden state for token $t$, from which the next token $y_{t}$ is sampled according to the decoding policy [brown2020language]. The Scheduler collects $y_{t}$ and feeds it back into the pipeline to form $x_{1:t+1}$, advancing the process to step t+1t{+}1. The loop continues until the model emits an end-of-sequence token $\langle\texttt{EOS}\rangle$ or another termination condition is met.
Task-Type Indistinguishability and Robustness. The communication protocol, message format, and API surface used by Inferencers during inference are identical to those used by Verifiers in the subsequent verification phase [yuan2022decentralized]. Consequently, no node can distinguish in real time whether it is performing live inference or participating in a verification probe. This indistinguishability eliminates the opportunity for strategic deviation and simplifies implementation symmetry across both phases. Combined with VRF-driven role assignment, signed commitments, and auditable relays, the inference scheme produces a verifiable, tamper-evident execution trace with negligible online overhead.
@@ 5.6 Verification Scheme
During autoregressive inference, an LLM executes two phases: a prefill pass—where the prompt tokens are processed in parallel to populate the key–value (KV) cache—and a decode phase—where tokens are generated sequentially with causal dependencies on all previous outputs [brown2020language, shoeybi2019megatron]. The decoding phase dominates latency and is inherently serial, leading to limited GPU utilization and memory-bound performance [rajbhandari2022deep]. However, verification differs fundamentally: once the full output sequence $y_{1:T}$ is known, the verifier is no longer constrained by sequential dependencies. Instead, it can recompute all token representations simultaneously under the same causal mask, yielding a single parallel pass equivalent to prefill over the concatenated prompt–output sequence [sun2024zkllm, zhang2020delphi].
@@
and let $\mathcal{S}=[S^{\text{prompt}}\|S^{1}\|\dots\|S^{T}]$ denote the concatenated hidden states across all positions. Given the causal structure of the Transformer, the verifier can execute a single forward pass to reconstruct all hidden states $\{S^{(t)}_{i}\}_{t=1}^{T}$ for every segment $i\in[1,L]$ without performing incremental token decoding [vaswani2017attention]. This transformation replaces $T$ serial decode iterations with a single batched forward computation, eliminating redundant KV-cache updates and reducing cross-node communication. In practice, this yields significantly improved GPU utilization, especially on architectures with high Tensor Core throughput [rajbhandari2022deep].
@@
yielding near-linear scaling in the number of participating verifier groups [rajbhandari2022deep, zhang2020delphi].
@@
Prefill Computation. After the inference phase concludes, the Scheduler initiates segment-local verification by distributing the required input states to each verifier [qu2025zkgpt]. For segment $i>1$, the Scheduler sends the sequence $\{S_{i}^{(t)}\}_{t=1}^{T}$—the hidden states produced by the upstream inferencer for all tokens. For the first segment ($i=1$), the Scheduler instead provides the full tokenized input $\texttt{prompt}\|y_{1:T}]$, corresponding to the raw embedding input of the model. Upon receipt, each verifier executes a full-sequence prefill limited to its hosted layers:
@@
Importantly, the communication interface, data serialization format, and RPC protocol used during verification are identical to those used in live inference [yuan2022decentralized]. This design choice enforces task-type indistinguishability: from the perspective of any node, it is impossible to discern whether the current request corresponds to live inference or a verification probe [bonneau2015sok]. Such indistinguishability eliminates mode-specific behavior and prevents nodes from selectively deviating or shortcutting computation when acting as verifiers. Upon completion, each verifier must return its computed outputs $\{S^{(t)}_{i+1}\}_{t=1}^{T}$ through the same interface as during inference, producing an auditable trace indistinguishable in structure and timing.
Commitment Submission. Once the local recomputation completes, the Scheduler transitions verifiers into the commit phase. Each verifier constructs a tamper-evident cryptographic commitment to its final outputs—specifically, the hidden state corresponding to the last token ($t=T$) at its segment’s output boundary. This commitment is instantiated as a Merkle root [merkle1989certified]:
@@
and publishes the commitment on-chain through the designated verification contract [chen2024zkml, sun2024zkllm].
This on-chain record provides a permanent anchor for later verification steps. In subsequent sampling or dispute phases, verifiers may be required to disclose specific slices or elements of the hidden state $\{S^{(T)}_{i+1}\}$, accompanied by Merkle inclusion proofs that can be efficiently checked against the posted root [bonneau2015sok, micali1999verifiable]. This approach ensures both compactness and auditability: the entire verification process involves only lightweight commitments and selective disclosures rather than full tensor uploads, keeping blockchain overhead minimal while preserving cryptographic integrity.
In summary, the verification scheme leverages the known output sequence and the pre-recorded intermediate states to convert sequential autoregressive computation into parallel, segment-local prefill verification. Combined with Merkle commitments and signature anchoring, this design achieves strong verifiability and high throughput, ensuring that even a single honest verifier can detect and prove any inference deviation with overwhelming probability [qu2025zkgpt, sun2024zkllm].